# Emergent ac Effect in Nonreciprocal Coupled Condensates

### v1

@@ Abstract
Figure 1: (a) Schematic of three mutually nonreciprocally coupled condensates, with asymmetric couplings $J_{L}$ and $J_{R}$. (b) Chiral ferromagnetic phase, where all condensates share the same phase and undergo collective precession at a constant frequency. (c) Chiral vortex phase, characterized by a phase winding of $2\pi$ across the three sites; relative phases are locked while the global phase rotates uniformly. (d) Spontaneous ac phase, featuring two emergent frequencies: one associated with global phase rotation and the other with oscillations of the relative phases. Dashed red: relative phase evolution $\theta_{ij}$; solid lines: individual phases $\theta_{i}(t)$ [19].
@@
Figure 3: (a) Steady-state oscillation amplitude of $\theta_{21}$ as a function of coupling $J$ for $D/\gamma=1.2$ and $G/\gamma=0.1$. (b) Projected trajectories of persistent oscillations on the reduced phase space $\mathcal{M}$ at $J=0$, shown on the $\mathbb{T}^{2}$ submanifold of relative phases, for varying values of $G$ (with $\gamma=3$, $D=5$). Inset: oscillation frequency as a function of $D$, for $G=0$ (red), $G=1$ (green), and $G=3$ (blue), with $\gamma=3$ fixed. Both plots are produced from long-time numerical solutions of the generalized Josephson equation, starting near the ferromagnetic fixed point [35].
@@
We show the phase diagram in Fig. 2(a) for $J=0$ (see SM for more cases [34]). Interestingly, the ferromagnetic phase does not immediately transition into a (anti)vortex state. Instead, it first enters a dynamical regime characterized by persistent oscillations on the reduced phase space $\mathcal{M}$, with no equilibrium analogue. When the dissipative coupling $G$ exceeds a critical threshold, the system eventually settles into an (anti)vortex state. Figure 2(b) shows the steady-state oscillation amplitude of $\theta_{21}$ for fixed $D$. For $G<0$, the system resides in a ferromagnetic state that breaks the global U(1) symmetry. This state is generically dynamical for finite $J$: the Goldstone mode $\Phi$ exhibits uniform rotation at frequency $\Omega=-2J$, dynamically restoring U(1) symmetry on average, a hallmark of the chiral phase in nonreciprocal systems [18]. However, unlike previous works that rely on stochastic noise to stabilize chiral phases [18], the interacting condensate system realizes them purely through its internal dynamical structure, without invoking any other ingredients. We show such chiral ferromagnetic rotation in Fig. 3(a) and also include an example of the emergence of this chiral ferromagnetic phase in the condensate dynamics in Supplemental Video 1 [34], starting from an initial state near the ferromagnetic fixed point.
@@
As $G$ becomes positive, the system enters a richer dynamical regime: beyond global phase rotation (with frequency proportional to $J$), the system develops autonomous oscillations on $\mathcal{M}$ at a distinct frequency. To visualize this, we project the oscillatory trajectory onto the $\mathbb{T}^{2}$ submanifold in Fig. 3(b). These persistent relative-phase oscillations generate oscillatory particle currents between condensates, accompanied by corresponding modulations in particle numbers. Interestingly, this ac-Josephson-like effect emerges intrinsically from the nonreciprocity and nonlinearity of the system. It is principally distinct from conventional ac effects driven by external bias, which is set to zero in this study. We show this ac phase in the condensate dynamics, with two emergent frequencies, in Supplemental Video 2 [34]. From the perspective of dynamical systems, the transition from the ferromagnetic state to persistent ac oscillations is a manifestation of a Poincaré–Andronov–Hopf bifurcation [37] occurring on the five-dimensional manifold $\mathcal{M}$, which originates from a mechanism distinct from that of the chiral phase. At the Hopf threshold $G_{c}=2\left[(3J^{2}+\gamma^{2})^{1/2}-\gamma\right]$, oscillations of relative angles emerge with amplitude $A$ scaling as $A^{2}\propto G/2-\gamma+\sqrt{(G+\gamma)^{2}-9J^{2}}$. In Fig. 3(b), we observe a clear increase in oscillation amplitude with increasing $G$; its dependence on $J$ is also displayed in Fig. 3(a). Meanwhile, the oscillation frequency is shown to be $\omega=\sqrt{3}D+\mathcal{O}(G-G_{c})$ [see inset of Fig. 3(b)]. The ac phase therefore exhibits spontaneous breaking of time-translation symmetry through the emergence of two independent frequencies, reminiscent of time-crystalline behavior [20]: one originates from the coherent rotation of the U(1) zero mode (with frequency $\Omega$) and the other from stable limit cycles on $\mathcal{M}$ (with frequency $\omega$).
@@
Figure 4: (a) Instability boundary of the vortex phase for different $J$. Red: $J=0$; orange: $J/\gamma=-\sqrt{3}$; green: $J/\gamma=\sqrt{3}$. The shaded region marks the parameter space where the vortex state is unstable for $J=0$. (b) Upper panel: In the nonreciprocal Kuramoto model [34], both the ferromagnetic state (green) and the vortex state (purple) lose stability exactly at $G=0$, switching directly into each other. Lower panel: For the full dynamics, the ferromagnetic phase (green) first destabilises at a positive $G$, entering an intermediate ac state (red) before reaching the vortex phase (purple); conversely, the vortex phase becomes unstable at a negative $G$. The separation of these thresholds evidences hysteresis. (c) Phase diagram showing the vortex phase and its transition to other phases for $J=0$, which is obtained numerically. The yellow curve shows the analytically obtained instability boundary given by Eq. (9), which agrees well with the numerics. (d) Frequency $\omega$ of the steady state as a function of $D$ for $J=0$ and $G/\gamma=-0.98$. (e) Limit‑cycle trajectories projected onto the relative‑phase torus $\mathbb{T}^{2}$ for $J=0$ and $G/\gamma=-0.98$. From inner to outer trajectories, we set $D/\gamma=2.7$, $2.6$, $1.5$. Inset: oscillation amplitude versus $D$. Plots (c)-(e) are obtained from numerical solutions of the generalized Josephson equations, starting near the vortex fixed point [38].
@@
We show this instability boundary in Fig. 4(a) for different exchange coupling $J$. We stress that this condition differs markedly from that of the ferromagnetic state (8), which is not surprising since the analysis is carried out around different fixed points. It is clear that in certain parameter region, both vortex and ferromagnetic phases are stable, which gives rise to the hysteresis discussed below. This is in sharp contrast to the nonreciprocal Kuramoto model [34], where the transition is symmetric and the instability is always triggered at $G=0$ from both directions as shown in the upper panel of Fig. 4(b). The different nonlinearity around different phases leads to hysteretic behavior: the onset of instability depends on the direction of the transition, shown in the lower panel of Fig. 4(b). In other words, both the ferromagnetic and (anti)vortex states can coexist as stable fixed points for the same parameters [34], and the selected state depends on the history (initial condition or sweep direction), giving rise to hysteretic phenomena.
We present the phase diagram in Fig. 4(c) for $J=0$, obtained by solving the dynamics numerically by starting from the vortex phase in contrast to Fig. 2(a) where we start from a ferromagnetic state. The yellow curve in Fig. 4(c) shows the analytically obtained instability boundary (9), which agrees well with the numerics. We note that the vortex state persists within a finite window of negative dissipative coupling, $G<0$. This stability, absent in the nonreciprocal Kuramoto model [34], stems from dynamical feedback: the always‑stable number sector ($\mathbb{R}_{+}^{3}$) couples to the phase sector ($\mathbb{T}^{2}$) via the antisymmetric exchange $D$, effectively renormalizing the phase dynamics and shifting the vortex instability to more negative $G$. A larger $D$ strengthens this feedback, further widening the stability window as observed in Fig. 4(c). This vortex phase is characterized by a rotating Goldstone mode $\Phi$ with frequency $\Omega=J+\sqrt{3}D$. As a result, time-translation symmetry is spontaneously broken, while the global U(1) symmetry is dynamically restored on average. We show the stabilization of this chiral vortex state, with an initial condition close to the vortex fixed point, in Supplemental Video 4 [34].
@@ Acknowledgements.
Acknowledgments. We thank Even Thingstad and Kouki Nakata for insightful discussions. This work was supported by the Georg H. Endress Foundation and by the Swiss National Science Foundation, NCCR SPIN (grant number 51NF40-180604).

### v2

@@ Abstract
Figure 1: (a) Schematic of three mutually nonreciprocally coupled condensates, with asymmetric couplings $J_{L}$ and $J_{R}$. (b) Chiral ferromagnetic phase, where all condensates share the same phase and undergo collective precession at a constant frequency. (c) Chiral vortex phase with a phase winding of $2\pi$; relative phases are locked while the global phase rotates uniformly. (d) Spontaneous ac phase, featuring two emergent frequencies: one associated with global phase rotation and the other with oscillations of relative phases. Dashed red: relative phase $\theta_{ij}$; solid lines: individual phases $\theta_{i}(t)$ [19].
@@
We show the phase diagram in Fig. 2(a) for $J=0$ (see SM for more cases [34]). Interestingly, the ferromagnetic phase does not immediately transition into a (anti)vortex state. Instead, it first enters a dynamical regime characterized by persistent oscillations on the reduced phase space $\mathcal{M}$, with no equilibrium analogue. When the dissipative coupling $G$ exceeds a critical threshold, the system eventually settles into an (anti)vortex state. Figure 2(b) shows the steady-state oscillation amplitude of $\theta_{21}$ for fixed $D$. For $G<0$, the system resides in a ferromagnetic state that breaks the global U(1) symmetry. This state is generically dynamical for finite $J$: the Goldstone mode $\Phi$ exhibits uniform rotation at frequency $\Omega=-2J$, dynamically restoring U(1) symmetry on average, a hallmark of the chiral phase in nonreciprocal systems [18]. However, unlike previous works that rely on stochastic noise to stabilize chiral phases [18], the interacting condensate system realizes them purely through its internal dynamical structure, without invoking any other ingredients. We show such chiral ferromagnetic rotation in Fig. 3(a) and also include an example in Supplemental Video 1 [34], starting from an initial state near the ferromagnetic fixed point.
@@
As $G$ becomes positive, the system enters a richer dynamical regime: beyond global phase rotation (with frequency proportional to $J$), the system develops autonomous oscillations on $\mathcal{M}$ at a distinct frequency. To visualize this, we project the oscillatory trajectory onto the $\mathbb{T}^{2}$ submanifold in Fig. 3(b). These persistent relative-phase oscillations generate oscillatory particle currents between condensates, accompanied by corresponding modulations in particle numbers. Interestingly, this ac-Josephson-like effect emerges intrinsically from the nonreciprocity and nonlinearity of the system. It is principally distinct from conventional ac effects driven by external bias, which is set to zero in this study. We show this ac phase, with two emergent frequencies, in Supplemental Video 2 [34]. From the perspective of dynamical systems, the transition from the ferromagnetic state to persistent ac oscillations is a manifestation of a Poincaré–Andronov–Hopf bifurcation [37] occurring on the five-dimensional manifold $\mathcal{M}$, which originates from a mechanism distinct from that of the chiral phase. At the Hopf threshold $G_{c}=2\left[(3J^{2}+\gamma^{2})^{1/2}-\gamma\right]$, oscillations of relative angles emerge with amplitude $A$ scaling as $A^{2}\propto G/2-\gamma+\sqrt{(G+\gamma)^{2}-9J^{2}}$ (for the coefficient of proportionality see SM [34]). In Fig. 3(b), we observe a clear increase in oscillation amplitude with increasing $G$; its dependence on $J$ is also displayed in Fig. 3(a). Meanwhile, the oscillation frequency is shown to be $\omega=\sqrt{3}D+\mathcal{O}(G-G_{c})$ [see inset of Fig. 3(b)]. The ac phase therefore exhibits spontaneous breaking of time-translation symmetry through the emergence of two independent frequencies, reminiscent of time-crystalline behavior [20]: one originates from the coherent rotation of the U(1) zero mode (with frequency $\Omega$) and the other from stable limit cycles on $\mathcal{M}$ (with frequency $\omega$).
@@
Figure 3: (a) Steady-state oscillation amplitude of $\theta_{21}$ as a function of coupling $J$ for $D/\gamma=1.2$ and $G/\gamma=0.1$. (b) Projected trajectories of persistent oscillations on the reduced phase space $\mathcal{M}$ at $J=0$, shown on the $\mathbb{T}^{2}$ submanifold of relative phases, for varying values of $G$ (with $\gamma=3$, $D=5$). Inset: oscillation frequency as a function of $D$, for $G=0$ (red), $G=1$ (green), and $G=3$ (blue), with $\gamma=3$ fixed [35].
Figure 4: (a) Instability boundary of the vortex phase for $J/\gamma=0,-\sqrt{3},\sqrt{3}$ (red, orange, and green curves). The shaded region is the parameter regime where the vortex state is unstable for $J=0$. (b)Upper panel: In the nonreciprocal Kuramoto model [34], the ferromagnetic (green) and vortex (purple) states lose stability simultaneously at $G=0$ and switch directly into each other. Lower panel: In the full dynamics, the ferromagnetic state destabilizes at positive $G$ into an intermediate ac state (red) before reaching the vortex phase, while the vortex state becomes unstable at negative $G$. (c) Phase diagram showing the vortex phase and its transition to other phases for $J=0$. The yellow curve shows the analytically obtained instability boundary given by Eq. (9). (d) Frequency $\omega$ of the steady state as a function of $D$ for $J=0$ and $G/\gamma=-0.98$. (e) Limit‑cycle trajectories projected onto the relative‑phase torus $\mathbb{T}^{2}$ for $J=0$ and $G/\gamma=-0.98$. From inner to outer trajectories, we set $D/\gamma=2.7$, $2.6$, $1.5$. Inset: oscillation amplitude versus $D$. Plots (c)-(e) are obtained from numerical solutions of the generalized Josephson equations, starting near the vortex fixed point [38].
@@
We show this instability boundary in Fig. 4(a) for different exchange coupling $J$. We stress that this condition differs markedly from that of the ferromagnetic state (8), since the analysis is carried out around different fixed points with different topology. It is clear that in certain parameter region, both vortex and ferromagnetic phases are stable [34], which gives rise to the hysteresis discussed below. This is in sharp contrast to the nonreciprocal Kuramoto model [34], where the transition is symmetric and the instability is always triggered at $G=0$ from both directions as shown in the upper panel of Fig. 4(b). The different nonlinearity around different phases leads to hysteretic behavior: the onset of instability depends on the direction of the transition, shown in the lower panel of Fig. 4(b). In other words, both the ferromagnetic and (anti)vortex states can coexist as stable fixed points for the same parameters [34], and the selected state depends on the history (initial condition or sweep direction), giving rise to hysteretic phenomena.
We present the phase diagram in Fig. 4(c) for $J=0$, obtained by solving the dynamics numerically by starting from the vortex phase in contrast to Fig. 2(a) where we start from a ferromagnetic state. The yellow curve in Fig. 4(c) shows the analytically obtained instability boundary (9), which agrees well with the numerics. We note that the vortex state persists within a finite window of negative dissipative coupling, $G<0$. This stability, absent in the nonreciprocal Kuramoto model [34], stems from dynamical feedback: the always‑stable number sector ($\mathbb{R}_{+}^{3}$) couples to the phase sector ($\mathbb{T}^{2}$) via the antisymmetric exchange $D$, effectively renormalizing the phase dynamics and shifting the vortex instability to more negative $G$. A larger $D$ strengthens this feedback, further widening the stability window as observed in Fig. 4(c). This vortex phase is characterized by a rotating Goldstone mode $\Phi$ with frequency $\Omega=J+\sqrt{3}D$. As a result, time-translation symmetry is spontaneously broken, while the global U(1) symmetry is dynamically restored on average. We show the stabilization of this chiral vortex state in Supplemental Video 4 [34].
@@ Acknowledgements.
Acknowledgments. We thank Even Thingstad and Kouki Nakata for insightful discussions. This work was supported by the Georg H. Endress Foundation, by the Swiss National Science Foundation, NCCR SPIN (grant no. 51NF40-180604), and by the Ibn Sina Global Scholarship. DL acknowledges the Deanship of Research at KFUPM and the Quantum Center for the support received under Grant no. CUP25102 and no. INQC2500, respectively.